OpenStudy (anonymous):

Given the equation Square root of 8x plus 1 = 5, solve for x and identify if it is an extraneous solution.

OpenStudy (solomonzelman):

\(\sqrt{8x+1}=5\) or \(\sqrt{8x}+1=5\) ?

OpenStudy (anonymous):

the first one :)

OpenStudy (solomonzelman):

raise both sides the the second power

OpenStudy (solomonzelman):

\(\large\color{red}{ \displaystyle \left(\color{black}{~\sqrt{8x+1}}~\right)^2\color{black}{=} \left(~\color{black}{5}~\right)^2 }\)

OpenStudy (anonymous):

okay so im super bad at math so hopefully this wont be wrong